OpenStudy (anonymous):

i need help with this one, can you please show me the steps.. i have a test tomorrow and i need to get it. thanks During a 4th of July weekend, 32 vehicles became trapped on the Sunshine Skyway Bridge while it was being repaved. A recent city ordinance decreed that only cars with 4 wheels and trucks with six wheels could be on the bridge at any given time. If there were 148 tires that needed to be replaced to due to damage, how many cars and trucks were involved in the incident?

OpenStudy (anonymous):

i have that x+y=32

OpenStudy (anonymous):

x+y=32 4x+6y=148

OpenStudy (anonymous):

and then that 4x+6y=148

OpenStudy (ash2326):

yeah.let x be no.of cars y be no. of trucks x+y=32 also no.of tyres of car 4x no. of tyres of trucks=6y 4x=6y=148

OpenStudy (anonymous):

solve for x in the first equation in terms of y, plug it into the x in second to solve for y, then plug that answer back into the first equation to solve for x. viola!

OpenStudy (anonymous):

how do i get how many cars and trucks were involved?

OpenStudy (anonymous):

okay, so x in terms of y would look like this: x+y=32 x=32-y plug this "x" into the second equation and solve for y 4(32-y)+6y=148 128-4y+6y=148 128+2y=148 2y=20 y=10 plug this "y" into the first equation and solve for x x+10=32 x=22

OpenStudy (anonymous):

22 cars, 10 trucks